Mathematics 241 – Multivariable Calculus Quiz 5 – October 11, 2013 Your Name: Let f(x, y) = e^{y} cos(x + y^{2}) for both parts below. A) (10) Compute \frac{∂f}{∂x}(x, y) and \frac{∂f}{∂y} (x, y).

B) (10) Find the directional derivative D_{(\sqrt{3}/2,−1/2)}f(π/2, 0).